OpenStudy (anonymous):

How do you solve a Proportion? EX. k/10 = 36/45

OpenStudy (callisto):

Solve k/10 = 36/45? Multiply both sides by 10..

OpenStudy (anonymous):

That doesn't make sense, I mean, I multiplied 36 and 45 by 10 but still, how do I figure out "k" then?

OpenStudy (callisto):

\[\frac{k}{10} = \frac{36}{45}\]Multiply 10 to both sides\[10(\frac{k}{10}) = 10(\frac{36}{45})\]So, you can isolate k on the left, ie you can solve k.

OpenStudy (anonymous):

Would the answer be 8? The stupid Sub. at school sucks and she texted the whole time and didn't even teach the lesson.

OpenStudy (callisto):

It is 8. I'm sorry to hear that :(